HETAV PATEL

Grade 11,

if 3 equel and same electric charges are forming an equilateral triangle, the whre will the potential be 0. (dont give the ans at the centre-thats wrong)

if 3 equel and same electric charges are forming an equilateral triangle, the whre will the potential be 0. (dont give the ans at the centre-thats wrong)

Grade:9

1 Answers

vikas askiitian expert
13 years ago

let three charges are placed at the vertices of equilateral triangle then net potential at any point is given by

V = V1 +V2 + V3

V = kq/r , so

V1,V2,V3 = Kq/r1 , kq/r2 , kq/r3

V = kq(1/r1+1/r2+1/r3)

since charges are same & of same sign so kq is taken out of brakett now

V is 0 if q is 0 or if r1,r2,r3 all are infinity

so , the only condition is that a point shoud be at infinity ...

inside the triangle only electric field can be 0 not potential...